OpenStudy (anonymous):

a tank contains 20kg of salt dissolved in 5000L of water.Brine contains 0.3kg of salt per L and enters the tank at 25L/min. How much salt is in the tank after 30 mins if it drains at 25L/min?

OpenStudy (anonymous):

@SolomonZelman help

OpenStudy (anonymous):

@dumbcow hlp

OpenStudy (dumbcow):

set up differential equation rate the salt changes = inflow of brine - outflow \[\frac{ds}{dt} = 25 (.3) - 25 (\frac{s}{5000})\] \[\frac{ds}{dt} = 7.5 - \frac{s}{200}\] \[\frac{ds}{dt} = \frac{1500 -s}{200} \] separate variables and integrate \[\int\limits \frac{ds}{1500-s} = \int\limits \frac{dt}{200}\] \[- \ln (1500-s) = \frac{t}{200} + C\] \[1500 - s = C e^{-t/200}\] \[s(t) = 1500 - C e^{-t/200}\] plug in initial value of s(0) = 20 C = 1480 \[s(t) = 1500 -1480e^{-t/200}\] Finally plug in t=30 to get answer